Field Notes
<--2025-02-18 10:10:00 | 2025-03-25 11:30:00-->Other sites visited today: Mayberry | East End | Rush Ranch | Bouldin Corn | Bouldin Alfalfa | Gilbert Tract
Phenocam link: Searching...
Hill Slough: 2025-03-07 09:35:00 (DOY 66)
Author: Daphne Szutu
Others: Arman
Summary: Breezy, replaced 7500 bird spike, SWout has NANs--moved to different channel, sonic rangefinder bad--needs replacement, new pantyhose sock for miniDOT
2025-03-07 Hill Slough Arman and I arrived at 9:35 PST. The water level was high and the staff gauge read 6.0ft at 9:57 PST. It was cool, sunny, and increasingly breezy. We dropped off one pre-cut plywood piece by the tower for future replacement of the lower boardwalk. I downloaded met, cam, and USB GHG data. Arman cleaned flux and rad sensors and topped off the wash reservoir. Arman also added a new bird spike for the 7500; we don’t know where the previous one disappeared to. SWout still had NANs even after I checked the wiring at the last site visit. Today I moved it to a new mux channel and sent a new datalogger program. The sonic rangefinder data seemed to go bad after our last site visit. Doesn't look tidal at all. We didn't mess with it last time, but timing is a weird coincidence. This is the sensor that got wet and corroded during king tides. Joe removed the corrosion and re-wired it, but we probably need to replace the sensor and raise its mount. Eventually the CR1000 program should be edited to remove the channel water sensors after they are moved to the second Hill Slough site on the east side of Grizzly Island Road. We could also remove the Decagon NDVI sensors here since we now use broadband NDVI. Arman measured the surface and porewater conductivity. Surface, 1974uS, 12.4C Porewater, 6.23mS, 13.7C The 7500 read: 440ppm CO2, 286 mol/m3 H2O, 14.1C, 102.2 kPa, 97 SS – forgot to write after cleaning The 7700 read: 2.08 ppm, RSSI 59 – 88 after cleaning At the channel, I downloaded miniDOT data, swapped its desiccant, and put on a fresh pantyhose sock. We left at 10:00 PST. |
0 photos found
4 sets found
21 HS_met graphs found
4 HS_flux graphs found
1 HS_cam graphs found
Can't check HS_Processed
2 HS_Processed graphs found